OpenStudy (anonymous):

Max and Maggie have to clean the house. It takes Max 12 hours to clean the house, while Maggie can complete the task in 4 hours. Their sister says that it will take 3 hours to complete if they work together. Explain each step in solving this equation, and determine if the sister is correct or not. @SolomonZelman

OpenStudy (anonymous):

@robtobey Help

OpenStudy (imstuck):

This is a work problem...I can probably help you...

OpenStudy (anonymous):

Alright thanks

OpenStudy (imstuck):

Ok, if Max can clean the whole house in 12 hours, he will get 1/12 of the job done in one hour, right?

OpenStudy (anonymous):

Alright

OpenStudy (imstuck):

Maggie can clean the house in 4 hours, she will get 1/4 of the job done in one hour. We need to base their work on hours, individually first, to see if the sister is right. The equation of them working together will look like this then:

OpenStudy (anonymous):

Alright I'm Starting to understand

OpenStudy (imstuck):

\[\frac{ 1 }{ 12 }+\frac{ 1 }{ 4 }=\frac{ 1 }{ t }\]You are solving for t, how long it will take them to do the job together.

OpenStudy (anonymous):

Nice

OpenStudy (imstuck):

That's Max's one hour of work plus Maggie's one hour of work. It has to equal the total timee to do the job. The sister says t = 3, so let's find out. We need to find the LCM of those denominators.

OpenStudy (anonymous):

1/3 =1/x ?

OpenStudy (anonymous):

X = 3

OpenStudy (imstuck):